OpenStudy (moonmama):

Can someone please show me how to do this? "Simplify each expression. Use positive exponents. m^3 n^(–6) p^0"

satellite73 (satellite73):

any number to the power of zero is one

satellite73 (satellite73):

and \[b^{-z}=\frac{1}{b^z}\] so \[n^{-6}=\frac{1}{n^6}\]